OpenStudy (anonymous):

Use integration by parts to prove the following reduction formulae:

OpenStudy (anonymous):

\[\int\limits_{}^{}x^{n}\ln(x) dx=\frac{ x^{n+1} }{ (n+1)^{2} }[[1+(n+1)\ln(x)]+c\]

OpenStudy (zepp):

Integration by part: \[\large \int udv = uv-\int vdu\]

OpenStudy (zepp):

Let u be \(\large x^n\) du=\(\large nx^{(n-1)}\) dv=ln(x)\[\large v=\int ln(x)dx=x(ln(x)-1)+c\]

OpenStudy (zepp):

Apply the integration by part formula:\[\large \int x^nln(x)dx=ln(x)x^n-\int (x(ln(x)-1)nx^{n-1}dx\]

OpenStudy (zepp):

Um, should use du = ln(x) instead..

OpenStudy (anonymous):

thats is what i did, i used u=lnx.

OpenStudy (zepp):

u=ln(x) du=1/x dx dv=x^n v=integral of x^n dx = \(\LARGE \frac{x^{n+1}}{n+1}\)

OpenStudy (anonymous):

yup that's what i had!

OpenStudy (anonymous):

im down to:

OpenStudy (zepp):

So what's wrong with it? :(

OpenStudy (anonymous):

\[\frac{ x^{n+1} }{ n+1 }lnx-\int\limits_{}^{}\frac{ x^{n-1} }{ n+1 }dx\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

???

OpenStudy (zepp):

You forgot something in your integral

OpenStudy (zepp):

du is 1/x dx, not only dx

OpenStudy (zepp):

so the integral should be like \[\large \int \frac{x^{n+1}}{x(n+1)}dx\]

OpenStudy (anonymous):

ok

OpenStudy (zepp):

and since 1/(n+1) is just a constant, move it out of the integral \[\large \frac{1}{n+1}\int \frac{x^{n+1}}{x}dx\]

OpenStudy (anonymous):

alrighty.

OpenStudy (zepp):

and that boils down to \[\large \frac{ln(x)x^{n+1}}{x^{n+1}}-\frac{1}{n+1}\int x^ndx\]

OpenStudy (zepp):

\[\large \frac{ln(x)x^{n+1}}{x^{n+1}}-\frac{1}{n+1}\frac{x^{n+1}}{n+1}+c\\\large \frac{ln(x)x^{n+1}}{x^{n+1}}-\frac{x^{n+1}}{(n+1)^2}+c\]

OpenStudy (zepp):

oups, the denominator of the first fraction is n+1, typo, my bad

OpenStudy (anonymous):

that doesnt look like what the answer says it is tho in the quesiton stated? :S

OpenStudy (zepp):

\[\large \frac{ln(x)x^{n+1}}{n+1}-\frac{x^{n+1}}{(n+1)^2}+c\]

OpenStudy (zepp):

It will :)

OpenStudy (anonymous):

is that the same thing?

OpenStudy (zepp):

They are, in the question, they just forged them into one single fraction

OpenStudy (anonymous):

okay thanks!

OpenStudy (zepp):

np